OpenStudy (anonymous):

find the product (5c+4)(5c-4)

OpenStudy (anonymous):

is it 25c^2-16?

OpenStudy (anonymous):

you would have to know that \[(a+b)(a-b)=a^2-b^2\] using that, \[(5c+4)(5c-4)=(5c)^2-(4)^2\]

OpenStudy (anonymous):

oh......you already found it!

OpenStudy (luigi0210):

Good job!

OpenStudy (anonymous):

you are right.

OpenStudy (anonymous):

Thank you. I thought my answer was wrong

OpenStudy (anonymous):

yw..(: